OpenStudy (anonymous):

Which of the following is the contrapositive of the conditional below? If 3 + 4 = 6, then 2 · 5 = 10.

OpenStudy (zzr0ck3r):

A implies B then the contrapositive is not B implies not A

OpenStudy (anonymous):

?

OpenStudy (anonymous):

If 2 · 5 = 10, then 3 + 4 = 6. If 3 + 4 ≠ 6, then 2 · 5 ≠ 10. If 2 · 5 ≠ 10, then 3 + 4 ≠ 6. If 3 + 4 ≠ 6, then 2 · 5 = 10.

OpenStudy (anonymous):

The third one is the correct one

OpenStudy (zzr0ck3r):

so you have \(3+4=6\) then \(2*5=10\) you need \(not (2*5=10)\) then \not(3+4=6\)

OpenStudy (zzr0ck3r):

@Chuck_O_Beast please read the rules about giving out answers, it helps no body

OpenStudy (anonymous):

Ok I apologize

OpenStudy (anonymous):

thank you!!!!!!!!!

OpenStudy (zzr0ck3r):

you should understand why before you move on @princess1298 but thats just by advice

OpenStudy (anonymous):

ok ill listen

OpenStudy (zzr0ck3r):

just read and see whats going on

OpenStudy (anonymous):

A contrapositive statement is a statement that is in the opposite order than the original and in opposite format

OpenStudy (zzr0ck3r):

well that is a bad definition, but it might help to understand it.

OpenStudy (anonymous):

o i got ya thank you for explaining

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

Ur welcome